What are (i) complements and (ii) supplements of the following angles a. 40 degrees b. 80 degrees c. 36 degrees
Answer by Theo(13342)

the sum of complementary angles is 90 degrees.

the sum of supplementary angles is 180 degrees.

complement of 40 is 50 because 40 + 50 = 90

supplement of 40 is 140 because 40 + 140 = 180

complement of 80 is 10 because 80 + 10 = 90

supplement of 80 is 110 because 80 + 110 = 180